ABC and CVP Analysis: Multiple Products

Good Scent, Inc., produces two colognes: Rose and Violet. Of the two, Rose is more popular. Data concerning the two products follow:

Rose Violet
Expected sales (in cases) 49,500 9,900
Selling price per case $99 $80
Direct labor hours 33,700 6,000
Machine hours 9,050 3,500
Receiving orders 50 24
Packing orders 103 51
Material cost per case $51 $41
Direct labor cost per case $10 $5

The company uses a conventional costing system and assigns overhead costs to products using direct labor hours. Annual overhead costs follow. They are classified as fixed or variable with respect to direct labor hours.

Fixed Variable
Direct labor benefits $          —   $178,650
Machine costs 219,500* 218,350
Receiving department 232,000  
Packing department 134,000  
   Total costs $585,500   $397,000

* All depreciation

Required:

1. Using the conventional approach, compute the number of cases of Rose and the number of cases of Violet that must be sold for the company to break even. In your computations, round variable unit cost to the nearest cent and round the number of break-even packages to the nearest whole number.

Break-even cases of Rose
Break-even cases of Violet

2. Using an activity-based approach, compute the number of cases of each product that must be sold for the company to break even. In your computations, round all computed amounts to the nearest cent and round the number of break-even packages to the nearest whole number.

Break-even cases of Rose
Break-even cases of Violet
Answer #1

Traditional = we would bifurcate overhead on the basis of labor hours

fixed variable
Overhead 585500 397000
direct labor hours [33700+6000] 39700 39700
per direct labor hour 14.75[585500/39700] 10

allocated to rose =

fixed cost = 585500/39700*33700=497011$

variable =10*33700=337000

allocated to violet = 585500/39700*6000=88489$

variable 10*6000= 60000$

Rose violet
Overhead allocated variable per case 6.81$[337000/49500cases] 6.06$[60000/9900]
direct material 51$ 41$
labor 10$ 5$
Total variable 67.81 52.06
sale 99$ 80$
contribution margin per case[sales-variable] 31.19$ 27.94$
fiexd cost 497011 88489
break even units [fixed cost/contribution margin per unit] 15935[497011/31.19$] 3167[88489/27.94$]

2.ABC COSTING uses specific activities to allocate cost

we will allocate cost on the following bases

direct labor benefit labor hours
machine cost machine hours
receiving department receiving orders
packing department packing orders
ROSE VIOLET
VARIABLE COST per unit
DIRECT LABOR BENEFIT 3.06[178650/39700*33700]/49500units 2.73[178650/39700*6000]/9900units
machine cost 3.18[218350/12550Total machine hours]*9050 / 49500units 6.15[218350/12550]*3500/9900
direct labor 10 5
direct material 51 41
Total variable cost 67.24 54.88
sales price 99 80
contribution margin [ sales- variable cost] 31.76 25.12
Total fixed cost
fixed cost machine 158285[219500/12550]*9050 61215[219500/12550]*3500
receiving department 156757[232000/74receiving orders]*50 75243[232000/74]*24
packing department 89623[134000/154]*103 44377[134000/154]*51
Total fixed cost 404665[158285+156757+89623] 180835[61215+75243+44377]
break even units 12741[404665/31.76] 7199[180835/25.12]

total machien hours =9050+3500=12550

total receiving orders= 50+24

total packing =103+51=154

> How did u get 14.75?

Jonathan Belmares Wed, Nov 17, 2021 5:00 PM